OpenStudy (anonymous):

Help solve? x^2+5x=-6 I got an answer, but I don't think it's right.

OpenStudy (anonymous):

(x+3)(x+2)=0

OpenStudy (anonymous):

x^2+5x+6=0 (x+3)(x+2)=0

OpenStudy (anonymous):

What did you get for your answer?

OpenStudy (anonymous):

-3 and -2

OpenStudy (anonymous):

yep thats right :)

OpenStudy (anonymous):

Thank you!